OpenStudy (anonymous):

4x + y = 11. x + 2y = 8

OpenStudy (anonymous):

what are you searching for

OpenStudy (arabpride):

combine like terms: 4x + y = 11. x + 2y = 8

OpenStudy (radar):

I am assuming that your post represents a system of equations. Two common ways to solve are: "Elimination" or "Substitution. I choose to use the "elimination" method. This method you eliminate one of the variables (either x or y) ending up with an equation with only one unknown. I also choose to eliminate y. To do this use employ the rules of algebra.

OpenStudy (radar):

I will first multiply the first equation by 2. Remember equals multiplied by equals will produce a product which is also equals. 2(4x + y = 11) = 8x + 2y = 22 I did this so that I now have two equations that have identical coefficients of y. because I want to eliminate the y. 8x + 2y = 22 I now want to subtract the second equation from this -(x + 2y = 8) ---------------- 7x =14 7x = 14 dividing both sides by 7 we solve for x, x = 2

OpenStudy (radar):

Now substitute this value of x (2) in one of the original equations and then solve for y. Note you could use either equation. 2 + 2y = 8 subtracting 2 from both sides 2y = 8 - 2 = 6 2y=6 y = 3

OpenStudy (radar):

A final step you should get used to doing is to verify these values. by replacing them in the original equations. Like so: 4x + y = 11 getting 4(2) + 3 = 11, 8 + 3 = 11 x + 2y = 8, 2 + 2(3) = 8, 2 + 6 = 8

OpenStudy (radar):

Checks out, Now I know I am correct.